OpenStudy (anonymous):

3/7 of the apples in a box are red apples. the rest are green apples. there are 24 green apples. how many apples are there altogether? "I'm lost" (please help :0)

OpenStudy (amistre64):

24 = 4/7 of a box divide each side by 4 6 = 1/7 of a box. now just put 3/7 together 18 = 3/7 18+24 = 42?

OpenStudy (anonymous):

4/7 of the apples=24 7/7 '' '' '' =7/4 X 24 =42

OpenStudy (amistre64):

gilberts was nice :)

OpenStudy (anonymous):

thank you